OpenStudy (anonymous):

Solve. x2 = 10x – 20

OpenStudy (anonymous):

x=5+-sqrt(5)

OpenStudy (anonymous):

\[x^2-10x=20\] \[(x-5)^2=20+25=45\] \[x-5=\pm\sqrt{45}=\pm3\sqrt{5}\] \[x=5\pm3\sqrt{5}\]